OpenStudy (anonymous):

Provide the solution to the equation 8 =√(2)^(4x + 8). Using complete sentences, explain the strategy used in solving this equation the way you did.

OpenStudy (kira_yamato):

8 =√(2)^(4x + 8) By factorizing 4x + 8, we've 4(x+2) By the law of indices, sqrt(2^4(x+2)) = 2^(2(x+2)) So 2^(2(x+2)) = 2^3 Since bases are the same, compare power 2(x+2) = 3 x + 2 = 1.5 x = -0.5